OpenStudy (anonymous):

Which of the following is the radical expression of 5/a^7 ? ^5 square root a^7 ^7 square root a^5 5a^7 7a^5

OpenStudy (radar):

The expression is not clear, can you use the draw command and write it please.

OpenStudy (radar):

\[\sqrt[5]{a ^{7}}\]Is that the expression?

OpenStudy (anonymous):

yes

OpenStudy (radar):

\[a ^{^{7/5}}=\sqrt[5]{a ^{5}a ^{2}}=a \sqrt[5]{a ^{2}}\]

OpenStudy (anonymous):

oh, so would it be the second one?

OpenStudy (radar):

Oh I see they did not ask to simplify, so I guess they are looking for \[a ^{7/5}\] is that one of your options?

OpenStudy (radar):

If it looks like that yes, but I can't tell because you have not used the equation editor, or drew (write) the expression.

OpenStudy (anonymous):

is it \[^{7}\sqrt{a}^{5}\]

OpenStudy (anonymous):

thats the second choice

OpenStudy (radar):

O.K, thats it then.

OpenStudy (radar):

Wait a minute, they have the exponents switched the have, it should look like\[a ^{7/5}\] My bad, assuming the original problem looked like my second post>

OpenStudy (radar):

Do they have one that looks like that?

OpenStudy (anonymous):

no :/

OpenStudy (radar):

\[\sqrt[5]{a ^{7}}\] If that is the original one then it would look like this\[a ^{7/5}\]

OpenStudy (anonymous):

ohhh so its the first one, correct?

OpenStudy (anonymous):

you were right, the options are just given in a different form

OpenStudy (anonymous):

i think..

OpenStudy (radar):

or maybe vice versa, if the problem is\[a ^{7/5}=\sqrt[5]{a ^{7}}\]

OpenStudy (anonymous):

ahh, so uh, it is the first?

OpenStudy (radar):